OpenStudy (anonymous):

9x - y = -2 is what in slope form? ?.?

OpenStudy (anonymous):

Point - slope form?

OpenStudy (anonymous):

like when i do it i get -y=-2-9x

OpenStudy (anonymous):

y=9x+2

OpenStudy (anonymous):

\[ y-y _{1}=m(x-x _{1})\] this is point slope form, equation of a line is y = mx+b